There are 42 male employees and 33 females employees in a company. Find the ratio of

(a) the number of male employees to the number of female employees,

(b) the number of female employees to the total number of employees.


1
Expert's answer
2022-05-15T10:51:38-0400

(a)


"\dfrac{42}{33}=\dfrac{14}{11}"

(b)


"\dfrac{33}{42+33}=\dfrac{33}{75}=\dfrac{11}{25}"